# Bosch Research North America & Bosch Center for AI https://zixunh.github.io/3d-geer

### v1

@@ 3.3. Efficient Ray-Particle Association via PBF
Figure 5. An Illustration of PBF-CSF Association. Left: Each PBF and CSF is parameterized by four spherical angles, with the PBF additionally defined by four planes tangent to a 3D Gaussian ellipsoid. Right: The intersection between the PBF and CSF, unfolded onto the BEAP imaging plane.
@@
To efficiently compute a PBF, we first define two spherical angles (See Fig. A.1 in Appendix A) in the camera space as:
(10) $\theta=\arctan\left(\frac{\mathbf{d}_{c,x}}{\mathbf{d}_{c,z}}\right),\quad\phi=\arctan\left(\frac{\mathbf{d}_{c,y}}{\mathbf{d}_{c,z}}\right),$
@@
where $(\mathbf{d}_{c,x},\mathbf{d}_{c,y},\mathbf{d}_{c,z})$ denotes the normalized direction vector pointing to the camera (view) space point $\mathbf{d_{c}}$. The angles $\theta$ and $\phi$ correspond to the ray’s incidence angles in the horizontal ($xz$) and vertical ($yz$) planes, respectively, each ranging from $-\frac{\pi}{2}$ to $\frac{\pi}{2}$.
We define a PBF by four planes — two aligned with the $\theta$ angle and two with the $\phi$ angle (See Figure 5):
(11) $\mathbf{g}_{\theta_{1,2}}=\left(-1,0,\tan\theta_{1,2},0\right)^{\top},\quad\mathbf{g}_{\phi_{1,2}}=\left(0,-1,\tan\phi_{1,2},0\right)^{\top},$
@@
Given a camera with extrinsic matrix $V=\left[R_{c}\mid\mathbf{t}_{c}\right]\in\mathrm{SE}(3)$ and a canonical-to-world transformation matrix $H$ (See Eq. 4), we apply the inverse transpose of the camera-to-canonical matrix $(VH)^{-1}$ to parameterize the corresponding canonical planes using the angular limits defined in view space. This transformation enables us to effectively derive the constraints for the PBFs. The canonical planes are given by:
(12) $\mathbf{g}_{u_{1,2}}$ $=(VH)^{\top}\mathbf{g}_{\theta_{1,2}}=-(VH)_{0}+\tan\theta_{1,2}(VH)_{2},$
$\mathbf{g}_{v_{1,2}}$ $=(VH)^{\top}\mathbf{g}_{\phi_{1,2}}=-(VH)_{1}+\tan\phi_{1,2}(VH)_{2},$
@@
(13) $\left((\lambda^{2},\lambda^{2},\lambda^{2},-1)\circ\mathbf{g}_{u}^{\top}\right)\cdot\mathbf{g}_{u}=0,$
where $\circ$ denotes the Hadamard (element-wise) product, and $\lambda$ is a parameter that controls the spread of the contour, typically set to $3$. An identical constraint is applied to $\mathbf{g}_{v}^{\top}$ in order to compute the corresponding angular bounds $\phi_{1,2}$.
The roots of these quadratic equations determine the tangent values of the angular bounds $[\theta_{1},\theta_{2}]$ and $[\phi_{1},\phi_{2}]$, which define the PBF angular limits. Notably, solving these quadratic equations under canonical space constraints respects the true 3D covariance and 3D mean of the Gaussian. This formulation is both efficient and eliminates the need for any intermediate approximation steps as adopted in EWA or UT (See Appendix E.1 for details).
3.4. Bipolar Equiangular Projection (BEAP)
We introduce a novel BEAP image representation for several reasons. First, it effectively represents images from large field-of-view (FoV) cameras without introducing FoV loss. Second, it allows image tiles and their corresponding Camera Sub-frustums (CSFs) to share an identical parameterization in terms of two $\theta$ and two $\phi$ angles, enabling highly efficient PBF association and optimal GPU parallelization. Third, the resulting ray sampling achieves a more uniform distribution in 3D space compared to conventional projective rendering, which empirically improves training efficiency and reconstruction quality, as shown in the ablation study in Sec. 5.2.
Figure 6. Comparison of Ray Distributions under Varying Projections. Each point represents a sampled ray, defined by the intersection between a ray from the optical center and the sphere surface. Our BEAP projection yields the most uniform ray distribution across the sphere surface.
@@ 3.4. Bipolar Equiangular Projection (BEAP)
(14) $(x,y)=\left\lfloor\frac{\mathrm{w}\theta}{\mathrm{FoV}_{x}}+\frac{\mathrm{w}+1}{2},\frac{\mathrm{h}\phi}{\mathrm{FoV}_{y}}+\frac{\mathrm{h}+1}{2}\right\rfloor,$
where $\mathrm{FoV}_{x},\mathrm{FoV}_{y}$ indicate the camera’s horizontal and vertical field-of-view. As illustrated in Figure 6, sampling rays uniformly in $(\theta,\phi)$ leads to more balanced spatial coverage within the view frustum. Our sampling strategy contrasts with projective sampling, which disproportionately allocates samples toward peripheral regions, and equidistant projection, which oversamples near the image center.
4. Experiments
Table 2. Quantitative Results on the ScanNet++ Dataset. Each method is trained using full FoV data but evaluated separately on the full FoV, central, and peripheral regions to assess performance under varying distortion levels. The region-wise metrics is weighted by pixel-wise masks, where LPIPS is omitted as it is not directly separable by region.
Method Full FoV Central Peripheral
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
FisheyeGS 27.81 0.946 0.139 32.44 0.956 23.28 0.914
EVER 29.47 0.924 0.167 29.93 0.924 28.72 0.925
3DGUT 30.64 0.944 0.150 31.87 0.945 28.84 0.937
3DGEER (Ours) 31.50 0.953 0.126 32.64 0.955 28.94 0.945
Table 3. Quantitative Results on the ZipNeRF Dataset. The experiments follow a cross-camera generalization setup, where each method is trained on either Fisheye (FE) or Pinhole (PH) data and evaluated separately on both Fisheye and Pinhole test sets.
@@ 4. Experiments
In this section, we evaluate our proposed method, 3DGEER, across multiple datasets featuring both pinhole and fisheye camera models.
Baselines. We compare 3DGEER with leading volumetric particle rendering methods targeting projective approximation errors in 3DGS (Kerbl et al., 2023). For large FoV fisheye inputs, we evaluate against FisheyeGS (Liao et al., 2024), EVER (Mai et al., 2024), and 3DGUT (Wu et al., 2025). Under pinhole settings, we additionally compare with 3DGS and 3DGRT (Moënne-Loccoz et al., 2024).
Datasets. To systematically evaluate performance under large FoV conditions, where projective approximation errors are most critical. We use five diverse scenes from the ScanNet++(Yeshwanth et al., 2023) dataset and the fisheye ZipNeRF(Barron et al., 2023), which has been largely overlooked in prior work. For conventional pinhole settings, we use the standard ZipNeRF and the widely adopted MipNeRF360 (Barron et al., 2021) datasets.
[…]
Implementation Details. To ensure fair comparison, we align our implementation with the original 3DGS in terms of initialization and hyperparameters on the ScanNet++ and MipNeRF360 datasets. For ZipNeRF, we adopt the training configurations of EVER and SMERF (Duckworth et al., 2024) for all candidate methods. Similar to most ray-based frameworks, our method does not directly access gradients of mean2D to control densification or pruning. Instead, we optimize using gradients of the view-space mean3D. Unless otherwise noted, all models are trained with consistent supervision strategies and architectural settings. For fisheye datasets (ScanNet++ and ZipNeRF), we precompute bijective grid mappings (Guo et al., 2025) and reproject all rendered outputs to the native KB-Fisheye image space (Kannala and Brandt, 2006) before evaluation. Experiments on ScanNet++ and MipNeRF360 are conducted using full-resolution images. For ZipNeRF, due to the high frame count and extremely large fisheye resolution, we use the maximum resolution (see Tab. 3) that allows all baselines to run reliably on a local desktop setup. Specifically, all experiments are conducted using an RTX 4090 GPU and 64GB of RAM as the default configuration.
Table 4. Quantitative Results on the MipNeRF360 Dataset. In addition to quality metrics, FPS is reported using an RTX 4090 as the default benchmark. Numbers sourced from original works are marked with (measured on RTX 6000 Ada) and (measured on RTX 5090).
@@
4.1. Evaluation on Large-FoV Novel View Synthesis
We evaluate 3DGEER against representative baselines on two fisheye datasets designed for benchmarking novel view synthesis. To ensure a more comprehensive evaluation, we incorporate additional analysis tailored to each dataset.
ScanNet++. As shown in Tab. 2, our method consistently outperforms all prior methods across all metrics, with clear margins. To further analyze each method’s performance under varying distortion, we separately evaluate reconstruction quality in the central and peripheral image regions (Detailed rule explained in Appendix G). As expected, FisheyeGS—while effective at enabling differentiable rendering for fisheye inputs—does not explicitly address projective approximation error. It performs well in the central region but degrades significantly in the periphery. Although all methods show some performance drop in peripheral areas, 3DGEER achieves the best results in both regions. Qualitatively, as illustrated in Fig. 7, 3DGEER delivers superior peripheral reconstruction compared to FisheyeGS, exhibits fewer artifacts and sharper details than EVER, and produces crisper object boundaries and textures—such as wires, keyboards, and wooden floors—when compared to 3DGUT.
@@ 4.1. Evaluation on Large-FoV Novel View Synthesis
ZipNeRF. Since the ZipNeRF dataset provides frame-wise aligned pinhole and fisheye datasets, we go beyond standard evaluation on the fisheye dataset by also conducting a cross-camera generalization evaluation—training on fisheye views and evaluating on pinhole, and vice versa—to further assess each method’s robustness and adaptability. As shown in Tab. 3, 3DGEER consistently outperforms all other methods across all metrics when trained on fisheye data, regardless of whether the test set is fisheye or pinhole. This highlights 3DGEER’s ability to fully leverage the advantages of large FoV training data through exact volumetric rendering. When training on pinhole data, 3DGEER also achieves significant PSNR improvements on both fisheye and pinhole test sets. In contrast, FisheyeGS performs well when both training and testing are done on pinhole images, but its performance drops sharply when evaluated on fisheye inputs. Similarly, 3DGUT exhibits a substantial decline when tested on fisheye views after training on narrower FoV data. Overall, EVER and 3DGEER demonstrate the most robust generalization across camera models—attributed to their preservation of geometric exactness—while 3DGEER consistently outperforms EVER across all scenarios. Qualitatively, as shown in Fig. 9, 3DGEER reconstructs finer details and exhibits greater robustness to artifacts than competing methods, both in challenging peripheral regions and cross-camera evaluation settings.
4.2. Evaluation on Normal-FoV Novel View Synthesis

### v2

@@ Discussion.
Figure 3: PBF-CSF Association. (Left) PBF defined by four planes tangent to a 3D Gaussian ellipsoid. (Right) The intersection between the PBF and CSF, unfolded onto the BEAP imaging plane.
3.2 Ray-Particle Association with Exact Projective Geometry
Given a ray-based Gaussian rendering method, both its final accuracy—particularly under large-FoV camera models—and its runtime efficiency depend heavily on the ray-particle association strategy, which has been largely overlooked in prior works. To determine which Gaussians contribute to each ray, we simplify the association problem by mapping Camera Sub-Frustums (CSFs) to Particle Bounding Frustums (PBFs)—the tight frustums enclosing each Gaussian at a certain distance from its center (Fig. 3). A Gaussian is assigned to a CSF if its PBF intersects the CSF. All Gaussians associated with a CSF are then considered contributors to every ray within that CSF.
@@ 3.2 Ray-Particle Association with Exact Projective Geometry
To efficiently compute a PBF, we first define two spherical angles (see Fig. E.1) in the camera space:
$\theta=\arctan\left(\frac{\mathbf{d}_{c,x}}{\mathbf{d}_{c,z}}\right),\quad\phi=\arctan\left(\frac{\mathbf{d}_{c,y}}{\mathbf{d}_{c,z}}\right),$ (6)
@@
where $(\mathbf{d}_{c,x},\mathbf{d}_{c,y},\mathbf{d}_{c,z})$ denotes the normalized direction vector pointing to the camera (view) space point $\mathbf{d_{c}}$. The angles $\theta$ and $\phi$ correspond to the ray’s incidence angles in the horizontal ($xz$) and vertical ($yz$) planes, respectively, each ranging from $\frac{\pi}{2}$ to $\frac{\pi}{2}$. We define a PBF by four planes — two aligned with the $\theta$ angle and two with the $\phi$ angle (see Fig. 3):
$\mathbf{g}_{\theta_{1,2}}=\left(-1,0,\tan\theta_{1,2},0\right)^{\top},\quad\mathbf{g}_{\phi_{1,2}}=\left(0,-1,\tan\phi_{1,2},0\right)^{\top},$ (7)
@@
Given a camera with extrinsic matrix $V=\left[R_{c}\mid\mathbf{t}_{c}\right]\in\mathrm{SE}(3)$ and a canonical-to-world transformation matrix $H$ (See Eq. B.4), we apply the inverse transpose of the camera-to-canonical matrix $(VH)^{-1}$ to parameterize the corresponding canonical planes using the angular limits defined in view space. This transformation enables us to effectively derive the constraints for the PBFs. Take the bounds of $\theta$ as an example, the canonical plane is given by:
$\mathbf{g}_{u_{1,2}}$ $=(VH)^{\top}\mathbf{g}_{\theta_{1,2}}=-(VH)_{0}+\tan\theta_{1,2}(VH)_{2},$ (8)
@@
$\left((\lambda^{2},\lambda^{2},\lambda^{2},-1)\circ\mathbf{g}_{u}^{\top}\right)\cdot\mathbf{g}_{u}=0,$ (9)
where $\circ$ denotes the Hadamard (element-wise) product, and $\lambda$ is a parameter that controls the spread of the contour, typically set to $3$. An identical constraint is applied to $\mathbf{g}_{v}^{\top}$ in order to compute the corresponding angular bounds $\phi_{1,2}$. The roots of these quadratic equations determine the tangent values of the angular bounds $\theta_{1},\theta_{2}]$ and $\phi_{1},\phi_{2}]$, which define the PBF angular limits. Notably, solving these quadratic equations under canonical space constraints respects the true 3D covariance $\Sigma_{c}$ and 3D mean $\bm{\mu}_{c}$ of the Gaussian in the camera space. Specifically, for the bounds of $\theta$, we have $c=\tan\theta_{1,2}$ satisfying:
$\mathcal{T}_{22}c^{2}-2\mathcal{T}_{02}c+\mathcal{T}_{00}=0,\quad\text{where}\quad\mathcal{T}_{ij}=\lambda^{2}\Sigma_{c}^{i,j}-\left(\bm{\mu}_{c}\bm{\mu}_{c}^{\top}\right)^{i,j}.$ (10)
This formulation is efficient to solve (See details in Sec. D.1) and eliminates the need for intermediate conic approximations, as in EWA or UT, as well as the screen-space constraints in HTGS and 2DGS that limit their FoV.
3.3 Bipolar Equiangular Projection (BEAP)
Figure 4: Comparison of Ray Distributions under Varying Projections. We project pixels from three imaging spaces onto the unit sphere to compare ray distributions. Among them, our BEAP projection achieves the most uniform coverage.
We introduce a novel BEAP image representation for several reasons. First, it effectively represents images from large-FoV cameras without introducing FoV loss. Second, it allows image tiles and their corresponding CSFs to share an identical parameterization in terms of two $\theta$ and two $\phi$ angles, enabling highly efficient PBF association and optimal GPU parallelization. Third, the resulting ray sampling achieves a more uniform distribution in the 3D space compared to conventional projective rendering, which empirically improves training efficiency and reconstruction quality.
@@ 3.3 Bipolar Equiangular Projection (BEAP)
$(x,y)=\left\lfloor\frac{\mathrm{w}\theta}{\mathrm{FoV}_{x}}+\frac{\mathrm{w}+1}{2},\frac{\mathrm{h}\phi}{\mathrm{FoV}_{y}}+\frac{\mathrm{h}+1}{2}\right\rfloor,$ (11)
where $\mathrm{FoV}_{x},\mathrm{FoV}_{y}$ indicate the camera’s horizontal and vertical field of view. As illustrated in Fig. 4, sampling rays uniformly in $(\theta,\phi)$ leads to more balanced spatial coverage within the view frustum. Our sampling strategy contrasts with projective sampling, which disproportionately allocates samples toward peripheral regions, and equidistant projection, which oversamples near the image center.
4 Experiments
In this section, we evaluate our proposed method, 3DGEER, across multiple datasets featuring both pinhole and fisheye camera models. All experiments are conducted using an RTX 4090 GPU and 64GB of RAM as the default configuration, unless otherwise stated. Full implementation details are provided in the Sec. G.
Datasets.
We evaluate performance on generic distorted cameras, where projective errors are critical, on all scenes from Fisheye ZipNeRF (Barron et al., 2023), five from ScanNet++ (Yeshwanth et al., 2023), and two from Aria (Lv et al., 2025). ZipNeRF/ScanNet++ cameras have 180 diagonal FoV, and Aria has 110 circular FoV. For pinhole settings, we use Pinhole ZipNeRF and MipNeRF360 (Barron et al., 2021).
Evaluation Metrics and Protocol.
[…]
Figure 6: Extreme-FoV Generalization. 3DGEER reconstructs complete scenes with fewer Gaussians, while FisheyeGS suffers from severe quality degradation.
Table 2: Comparison on the ScanNet++ Dataset. Each method is trained either on the full FoV or only the central region, and then evaluated on the full, central, and peripheral regions to assess robustness under varying distortions. Region-wise metrics are weighted by pixel masks, while LPIPS is excluded since it is not directly separable by region (Tab. K.2 shows full stats).
Method Training Training Full FoV Central Peripheral
FoV Space PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
3DGS Central Persp. - - - 31.26 0.948 - -
FisheyeGS Full Eq. 27.81 0.946 0.139 32.44 0.956 23.28 0.914
EVER Full KB 29.47 0.924 0.167 29.93 0.924 28.72 0.925
3DGUT Full Eq. 30.64 0.944 0.150 31.87 0.945 28.84 0.937
3DGEER (Ours) Central BEAP 29.93 0.949 0.130 32.86 0.954 26.64 0.930
3DGEER (Ours) Full BEAP 31.50 0.953 0.126 32.64 0.955 28.94 0.945
Table 3: Comparison on the ZipNeRF Dataset. The experiments follow a cross-camera generalization setup, where each method is trained on either Fisheye (FE, $1/8$ resolution) or Pinhole (PH, $1/4$ resolution) data and evaluated separately on both Fisheye and Pinhole test sets (Tab. K.9 shows full stats).
@@ 4 Experiments
4.1 Scaling Gaussians vs. Projective Exactness in Wide-FoV Cameras
(Celarek et al., 2025) argues that simply scaling the number of Gaussians (up to 1M) suffices to mitigate the linear approximation gap between splat-based and ray-based methods. However, this strategy has only been demonstrated on small object-centric scenes and does not address the challenges posed by wide-FoV distortions. To further investigate, we scale up the splat-based fisheye method (FisheyeGS) on ScanNet++ scenes by lowering its densification threshold, producing 3.6–5M Gaussians (H200).
As shown in Fig. 5, the red curve (FisheyeGS) and blue curve (3DGEER) show a clear PSNR gap under the same Gaussian count, implying an approximation gap in projective exactness (from 4k to 30k iterations, after which densification typically saturates). The light-blue curve further shows the scaled-up FisheyeGS, despite its substantial increase in memory and computation, performance saturates at 29.3 PSNR, whereas ours achieves 32.1 with only 550-700K Gaussians. Full statistics are reported in Tab. K.3.
4.2 Comprehensive Evaluation on All-FoV Novel View Synthesis
We first evaluate 3DGEER against representative baselines on both fisheye and pinhole benchmarks. To ensure comprehensive evaluations, we incorporate additional analysis tailored to each dataset.
ScanNet++. As shown in Tab. 2, 3DGEER consistently outperforms all baselines across metrics. To examine distortion effects, we separately evaluate central and peripheral regions (details in Sec. G). Competing methods perform reasonably at the center but degrade at the periphery, whereas 3DGEER maintains high fidelity throughout, producing sharper boundaries and fewer artifacts (Fig. K.4). Notably, even when trained only on central crops, 3DGEER still surpasses FisheyeGS across all regions.
Qualitatively (Fig. 6), we evaluate extreme-FoV generalization by training on the original dataset FoV ($180^{\circ}$ diagonal FoV) but testing at much wider FoVs ($180^{\circ}$ FoV, circular shape). 3DGEER reconstructs complete scenes with fewer Gaussians, while splatting-based methods such as FisheyeGS degrade significantly at the periphery (see more visuals in Fig. K.1-K.2 with extreme FoV).
Table 4: Quantitative Results on the MipNeRF360 Dataset. In addition to quality metrics, FPS is reported using an RTX 4090 as the default benchmark. Numbers sourced from original works are marked with (measured on RTX 6000 Ada) and (measured on RTX 5090).
@@ 4.2 Comprehensive Evaluation on All-FoV Novel View Synthesis
ZipNeRF. The dataset provides frame-wise aligned pinhole and fisheye views. Beyond standard fisheye evaluation, we conduct cross-camera generalization—training on fisheye and testing on pinhole, and vice versa—to assess robustness. As shown in Tab. 3, 3DGEER consistently outperforms all methods across metrics, fully exploiting large-FoV training data via geometrically exact rendering. Trained on pinhole, 3DGEER still yields notable PSNR gains on both views. In contrast, FisheyeGS performs well only on pinhole-to-pinhole, and 3DGUT drops sharply when tested on fisheye after pinhole training. Overall, EVER and 3DGEER generalize better due to geometric exactness, while 3DGEER is superior in all cases. Qualitatively (Fig. 7), 3DGEER recovers finer details and shows stronger resistance to artifacts in the periphery in the most challenging pinhole-to-fisheye setting.
Figure 7: Qualitative Results on the ZipNeRF Dataset with Cross-Camera Generalization. Performance differences are pronounced in the most challenging Pinhole (PH) train – Fisheye (FE) test setup, where our method demonstrates a significant advantage.
Table 5: Training-Time Transmittance Replacement on ScanNet++. Using splatting-based transmittance increases the Gaussian count due to larger approximation error, yet the perceptual gap remains compared with our projective-exact formulation (Tab. K.7 shows the full stats).
Transmittance Full FoV Central Peripheral Avg. Gaussian
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ Number (k) $\downarrow$
3DGS Splats 22.86 0.878 0.177 27.16 0.884 18.56 0.855 1396.4
FisheyeGS Splats 27.90 0.948 0.141 32.49 0.957 23.36 0.917 920.6
Projective-Exact (Ours) 31.50 0.953 0.126 32.64 0.955 28.94 0.945 591.8